Excel Online Structured Activity: Project risk analysis The Butler-Perkins Company (BPC) must decide between two mutually exclusive projects. Each costs $6,500 and has an expected life of 3 years. Annual project cash flows begin 1 year after the initial investment and are subject to the following probability distributions: Project A Probability Project B Cash Flows $6,000 $6,500 $7,000 Probability Cash Flows 0.2 0.6 0.2 0.2 0.6 0.2 $0 $6,500 $18,000 BPC has decided to evaluate the riskier project at 12% and the less-risky project at 10%. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the questions below What is each projects expected annual cash flow? Round your answers to two decimal places. Project A: Project B: $
Answer #1
Project A Project B
Probability Cash Flows Probable cash flows Probability Cash Flows Probable cash flows
0.2 6000 1200 0.2 0 0
0.6 6500 3900 0.6 6500 3900
0.2 7000 1400 0.2 18000 3600
NPV= 6500 NPV= 7500

?A= [(0.2)(6000-6500)2 + (0.6)(6500-6500)2+ (0.2)(7000-6500)2]= 500

CVA = 500/6500=0.076

BPC should choose Project B because it has a higher NPV compared to A

If Project B being negatively correlated be known, then Project B would be less risky. This would lead to the acceptance of Project B. if it is negatively correlated it would indicate that it would be profitable when the economy is down, thus means less risky and acceptance of the said project to reinforced